Suppose that high temperatures in College Place during the month of January have a mean of 37∘F. If you are told that Chebyshev's inequality says at most 6.6% of the days will have a high of 42.5∘F or more, what is the standard deviation of the high temperature in College Place during the month of January? Round your answer to one decimal place.
Answer #1

from Chebyshev's inequality % of K std deviation above values from mean =(0.5/k2)*100 =6.6

k =sqrt(50/6.6)=2.75

hence (42.5-37)/std deviation =k=2.75

std deviation =5.5/2.75 =2